# Trainable and Explainable Simplicial Map Neural Networks

### 11 Mar 2024 v2

@@ 6 Discussion and Limitations
Let us remark that SMNNs are in between an instance-based method and a multilayer perceptron. The previous definition of SMNNs ([4, 5]) shared advantages with instance-based methods such as the k-Nearest Neighbour Algorithm. Some of the advantages are: there was no training phase, it handles easily complex decision boundaries by barycentric subdivisions, it is effective in low-dimensional spaces, it adapts well to local patterns in the data, and the decision-making is transparent and interpretable. However, it was computationally expensive for large high-dimensional datasets due to the Delaunay triangulation computation and the required convex polytope, it suffered for overfitting and lacked generalization ability.. The proposed update in this paper provides a substitute for the convex polytope which reduces the number of points needed to compute the Delaunay triangulation. Delaunay computation is also less expensive thanks to the use of a support set $U$.
@@ 7 Experiments
First, the SMNN were trained using the gradient descent algorithm and the cross-entropy loss function during $500$ epochs for $n=2,3,4,5$ and $\kappa=1000,100,50,10$. Besides, the two-hidden-layer feed-forward neural network FFNN was trained using the Adam training algorithm [17] for $n=2,3,4,5$. Both training processes were carried out for $500$ epochs. The accuracy and loss results provided in Table 1 are the mean of $100$ repetitions. We can observe that both the SMNN and the FFNN have similar performance, but the SMNN generally reach lower loss values. The variance in the results was on the order of $10^{-8}$ to $10^{-5}$ in the case of the SMNN and of $10^{-5}$ to $10^{-2}$ in the case of the FFNN.
Third experiment: we compare the performance of the SMNN depending on the choice of the support set $U$. To do this, we applied two different methods to choose $U$. On the one hand, we use the $\varepsilon$-representative sets previously computed in the first experiment for different values of $\kappa$. On the other hand, we use the two outlier-robust subsampling methods presented in [18] for different values of Topological Density (TD). In Figure 8, examples of different support sets computed using the three different approaches are shown. Let us remark that the outlier-robust subsampling method can be tuned to keep outliers, which the authors call vital landmark set, or not, obtaining a representative landmark set. The two methods were tested on synthetic datasets composed of $1000$ points. The SMNN was trained for $1000$ epochs using the $75\%$ of the dataset and tested in the remaining $25\%$. The number of features $n$ of each dataset considered, the size $m$ of each support set computed, and the mean accuracy and loss results of $5$ repetitions when training the SMNN are provided in Table 2. In Table 3, the time for the computation of the barycentric coordinates is provided. As we can see, time of execution does not have to be directly related to the size of the support set and it may increase if many of the points to be evaluated are outside the Delaunay triangulation of $U$. In this experiment, the results suggest that $\varepsilon$-representative datasets provide better results than the other support sets. However, thorough theoretical studies should be developed to confirm the last statement. Such a study is outside of the scope of this paper.
@@ 8 Conclusions
The ideas presented in this paper can be extended in many different ways. In future work, we pretend to study less-data dependant approaches so that the Delaunay triangulation is needless. Besides, this architecture should be extended to Deep Learning models so that it can be applied to more complex classification problems such as image classification and provides extra explainability to Deep Learning models.

### 12 Mar 2024 v3

@@ 6 Discussion and Limitations
Let us remark that SMNNs are in between an instance-based method and a multilayer perceptron. The previous definition of SMNNs ([4, 5]) shared advantages with instance-based methods such as the k-Nearest Neighbour Algorithm. Some of the advantages are: there was no training phase, it handles easily complex decision boundaries by barycentric subdivisions, it is effective in low-dimensional spaces, it adapts well to local patterns in the data, and the decision-making is transparent and interpretable. However, it was computationally expensive for large high-dimensional datasets due to the Delaunay triangulation computation and the required convex polytope, it suffered from overfitting and lacked generalization ability. The proposed update in this paper provides a substitute for the convex polytope which reduces the number of points needed to compute the Delaunay triangulation. Delaunay computation is also less expensive thanks to the use of a support set $U$.
@@ 7 Experiments
First, the SMNN was trained using the gradient descent algorithm and the cross-entropy loss function during $500$ epochs for $n=2,3,4,5$ and $\kappa=1000,100,50,10$. Besides, the two-hidden-layer feed-forward neural network FFNN was trained using the Adam training algorithm [18] for $n=2,3,4,5$. Both training processes were carried out for $500$ epochs. The accuracy and loss results provided in Table 1 are the mean of $100$ repetitions. We can observe that both the SMNN and the FFNN have similar performance, but the SMNN generally reaches lower loss values. The variance in the results was on the order of $10^{-8}$ to $10^{-5}$ in the case of the SMNN and of $10^{-5}$ to $10^{-2}$ in the case of the FFNN.
Third experiment: we compare the performance of the SMNN depending on the choice of the support set $U$. To do this, we applied two different methods to choose $U$. On the one hand, we use the $\varepsilon$-representative sets previously computed in the first experiment for different values of $\kappa$. On the other hand, we use the two outlier-robust subsampling methods presented in [19] for different values of Topological Density (TD). In Figure 8, examples of different support sets computed using the three different approaches are shown. Let us remark that the outlier-robust subsampling method can be tuned to keep outliers, which the authors call vital landmark set, or not, obtaining a representative landmark set. The two methods were tested on synthetic datasets composed of $1000$ points. The SMNN was trained for $1000$ epochs using the $75\%$ of the dataset and tested in the remaining $25\%$. The number of features $n$ of each dataset considered, the size $m$ of each support set computed, and the mean accuracy and loss results of $5$ repetitions when training the SMNN are provided in Table 2. In Table 3, the time for the computation of the barycentric coordinates is provided. As we can see, the time of execution does not have to be directly related to the size of the support set and it may increase if many of the points to be evaluated are outside the Delaunay triangulation of $U$. In this experiment, the results suggest that $\varepsilon$-representative datasets provide better results than the other support sets. However, thorough theoretical studies should be developed to confirm the last statement. Such a study is outside of the scope of this paper.
@@ 8 Conclusions
The ideas presented in this paper can be extended in many different ways. In future work, we intend to study less data-dependent approaches so that the Delaunay triangulation is needless. Besides, this architecture should be extended to Deep Learning models so that it can be applied to more complex classification problems such as image classification and provides extra explainability to Deep Learning models.